OpenStudy (anonymous):

Add the complex expressions, and simplify. (9 + 9.4i) + (-8.6 – 4i)

OpenStudy (akshay_budhkar):

Add the imaginary part together and the real part together

OpenStudy (anonymous):

combine like terms

OpenStudy (akshay_budhkar):

basic rule is\[(c+di)+(a+bi)= (a+c)+(b+d)i\]

OpenStudy (anonymous):

ok i am lost

OpenStudy (akshay_budhkar):

just combine like terms together like in this question you have \[(9 + 9.4i) + (-8.6 – 4i)= (9-8.6)+(9.4-4)i\]

OpenStudy (anonymous):

ok so would you add the same ones together or would it be like 9 ^2

OpenStudy (mertsj):

Add the real parts and add the imaginary parts. what is 9+-8.6?

OpenStudy (akshay_budhkar):

you would add the same ones together

OpenStudy (anonymous):

0.4

OpenStudy (mertsj):

What is 9.4i -4i?

OpenStudy (mertsj):

Yes. so the real part is .4

OpenStudy (mertsj):

Now do the imaginary part

OpenStudy (akshay_budhkar):

you just need to separate the real and the imaginary parts for real you have 9-8.6=0.4 for imaginary (9.4-4)i = 5.4i

OpenStudy (mertsj):

9.4i-4i

OpenStudy (mertsj):

So the answer is .4+5.4i